# FastVGGT: Training-Free Acceleration of Visual Geometry Transformer

### v1

@@ 1 Introduction
Inferring the 3D geometric structure of a scene from visual inputs, is critical for enabling machines to understand and interact with the physical world. Recent advances in deep learning have catalyzed a paradigm shift in 3D geometric estimation [27, 32, 15, 25, 16], enabling a move from iterative, optimization-based pipelines to end-to-end neural networks that directly infer geometry from raw visual inputs. This transformation is exemplified by large-scale architectures like DUSt3R [27] and its follow-ups [30, 23, 11], which showcase a remarkable capacity to reason about complex geometric relationships across image pairs.
Building upon this line of research, VGGT [25] marks a significant advance. Its transformer-based, feed-forward architecture directly regresses key 3D attributes, including camera parameters, depth maps, and point tracks, to achieve highly stable and accurate reconstructions. While this establishes VGGT as a state-of-the-art framework for 3D scene understanding, its scalability is impeded by two critical bottlenecks. First, the model’s reliance on dense global token interactions across views or frames results in prohibitive computational costs. Although used techniques like Flash-Attention mitigate the memory complexity from $O(n^{2})$ to $O(nd)$, the underlying time complexity remains quadratic at $O(n^{2}d)$. Second, the global attention mechanism, essential for capturing inter-frame relations, is susceptible to error accumulation. As the token space expands with each new frame, minor inaccuracies are amplified, leading to significant prediction drift. Collectively, these limitations restrict VGGT’s applicability in large-scale scenarios and motivate the development of more efficient and scalable architectures.
@@
Motivated by the observation of attention redundancy, we adapt the training-free technique of token merging [2, 1, 24, 10] to enhance VGGT’s inference efficiency. Token merging consolidates redundant representations by partitioning tokens into source (src) and destination (dst) sets and merging each src token into its most similar dst counterpart. While effective in 2D vision tasks, its extension to structures designed for 3D geometry understanding remains underexplored. Unlike 2D settings that process single images, VGGT relies on cross-image correspondences, making direct application of token merging highly challenging. To address this, we introduce FastVGGT, a novel, training-free framework that strategically applies token merging to mitigate the Global Attention bottleneck. Our approach begins by preserving the foundational coordinate system. Specifically, tokens from the initial frame, which serves as the global reference for the entire scene, are designated as high-priority destination (dst) tokens and are exempt from being merged to ensure reconstruction stability. Furthermore, to maintain global consistency and preserve fine-grained details, we identify and retain the most salient tokens across all frames, allowing them to bypass the merging process entirely and participate directly in the attention computation. Finally, drawing inspiration from ToMeSD [1], we implement region-based random sampling within each subsequent frame. This ensures a spatially balanced selection of src and dst tokens, preventing critical information loss in localized regions during consolidation.
@@ 2.1 Feed-Forward 3D Reconstruction
Building on the foundations of traditional 3D reconstruction [13, 19, 28, 21, 12], recent end-to-end learning-based methods leverage neural networks to encode scene priors, substantially enhancing robustness and cross-dataset generalization. Early progress was marked by DUSt3R [27], which directly regresses view-consistent 3D point maps from only two RGB images without requiring camera calibration. Its successor, MASt3R [11], introduces confidence-weighted losses to approximate metric scale, further improving reconstruction quality. The current state-of-the-art, VGGT [25] scales this philosophy to a 1.2B-parameter transformer that jointly predicts camera intrinsics, extrinsics, dense depth, point maps, and 2D tracks. However, as input sequences grow longer, the global attention mechanism must capture inter-frame relations within an expanding token space. This not only increases computational overhead but also amplifies noise propagation, making long-sequence predictions more prone to drift. VGGT-Long [7] addresses the drifting issue by aligning sub-maps to suppress error accumulation, but at the cost of significantly reduced inference speed, undermining the efficiency of feed-forward 3D reconstruction. To overcome these challenges, we propose FastVGGT, which accelerates inference and mitigates error accumulation by reducing the number of tokens processed in Global Attention, thereby achieving a balance between efficiency and accuracy.
@@ 2.2 Token Merging
Visual token merging [2, 17, 31, 9] was first introduced as a training-free approach to improve the throughput of Vision Transformers (ViTs) [8]. It was later extended to reduce computational cost in tasks such as diffusion, video and language understanding, and video editing [3, 5, 20]. The method partitions tokens into two sets, matches each token with its most similar counterpart, merges them via average pooling, and concatenates the results. Although simple and effective, most applications remain limited to the image domain, while long-form video remains underexplored despite its spatiotemporal tokens being both redundant and interdependent. In parallel, other approaches to token reduction have emerged. TokenLearner [18] adaptively selects informative tokens with an MLP. PuMer [3] integrates pruning and merging for vision–language models. Pooling-based methods accelerate attention by averaging token embeddings. While sharing the goal of efficient transformers, our work targets the unique demands of feed-forward 3D reconstruction, where tokens must retain spatial precision and temporal coherence. To this end, we propose FastVGGT, which adapts token merging to VGGT, achieving substantial acceleration while maintaining reconstruction fidelity.
@@ 3.1 Visualization Findings
The phenomenon of attention similarity, often referred to as feature degradation, has been consistently reported in the DINO series [14, 22, 4]. In DINO, while the CLS token becomes increasingly discriminative, patch-level features gradually lose local consistency as they converge toward the CLS token, which undermines performance on dense prediction tasks. A similar limitation appears in VGGT: its global self-attention layers aggregate information across all tokens via weighted averaging, and without explicit regularization or task-specific constraints, the representational space is repeatedly compressed. This compression drives tokens to collapse along a dominant direction, eroding their distinctiveness. At a fundamental level, both DINO and VGGT lack explicit mechanisms to preserve token diversity, leading to a progressive loss of local variation.
It is important to note that the consequences of feature degradation differ markedly between DINO and VGGT. In DINO, the network must handle not only image classification but also dense prediction tasks such as segmentation, where local feature integrity is critical. As training progresses, patch tokens collapse toward the CLS token, weakening dense feature discrimination and directly impairing downstream performance. This motivates corrective strategies such as Gram Anchoring [22]. By contrast, VGGT follows a two-stage design: Global Attention and Frame Attention. Global Attention enforces global consistency by capturing holistic spatial–temporal relationships. Here, the apparent degradation—tokens collapsing toward a dominant subspace—can be interpreted not as a flaw but as deliberate distillation of global semantics. Frame Attention then reintroduces local variability, striking a balance between global abstraction and local differentiation. Moreover, the strong feature similarity observed in Global Attention exposes redundancy that can be exploited to reduce computational cost without compromising performance, providing a natural path to address VGGT’s speed bottleneck.
@@
Figure 4: Visualizations of the Global Attention maps in VGGT, using six representative tokens (including the camera token and several image tokens), show that at every stage the attention patterns of different tokens exhibit a strong degree of similarity. G-Attn and F-Attn denote Global Attention and Frame Attention, respectively.
@@ 3.2 Token Partitioning
Uniform Token Sampling. Finally, recognizing the dense prediction nature of 3D reconstruction, we ensure uniform intra-frame sampling to avoid local over-compression or redundancy. To this end, we assign dst and src tokens within each frame using a region-based random sampling strategy inspired by the success of ToMeSD [1] in diffusion models. Concretely, we first partition the input tokens by frame and arrange them into a 2D grid of image patches. Within each grid cell, dst tokens are sampled according to a predefined merging ratio with stride $K$, while the remaining tokens are designated as src tokens. This region-based strategy ensures spatially balanced merging and prevents artifacts such as the disappearance of large areas. Consequently, the merging process is more coherent, and the reconstructed scene better preserves global structural stability.
@@ 3.4 Token Unmerging Procedure
Dense 3D reconstruction requires per-token outputs (e.g., depth predictions). To satisfy this requirement, we adopt an unmerging operation, inspired by ToMeSD [1], which restores the original token resolution and maintains full compatibility with the VGGT architecture. Specifically, suppose two tokens $x_{1},x_{2}\in\mathbb{R}^{c}$ are merged into a single representation:
@@ 4.1 Experimental Setup
We evaluate FastVGGT on three benchmark datasets: ScanNet, NRGBD, and 7 Scenes. For ScanNet, which contains 1,500 scenes, we uniformly sample 50 scenes to form a reproducible benchmark, denoted as ScanNet-50. Our experiments focus on two tasks: camera pose estimation and point map reconstruction. Across both tasks, FastVGGT consistently achieves substantial speedups while maintaining accuracy. The overall architecture follows VGGT, comprising L=24 frame and global attention layers, with Flash-Attention2 [6] integrated to further accelerate inference. All experiments are conducted on a workstation with an NVIDIA A800 GPU (80 GiB VRAM).
@@
$\pi^{3}$ [29] OOM OOM OOM OOM OOM OOM OOM OOM
StreamVGGT[33] OOM OOM OOM OOM OOM OOM OOM OOM
Fast3R [30] 0.684 397.8s 0.701 97.3s 0.711 34.9s 0.723 4.8s
CUT3R [26] 0.786 34.8s 0.774 18.8s 0.775 11.1s 0.767 3.6s
VGGT [25] 0.471 724.6s 0.420 177.5s 0.416 131.4s 0.423 9.1s
@@
$\pi^{3}$ [29] OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM
StreamVGGT [33] OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM
Fast3R [30] 0.045 0.027 0.047 0.010 0.616 0.627 43.7s 0.040 0.021 0.056 0.013 0.639 0.657 5.5s
CUT3R [26] 0.179 0.121 0.097 0.043 0.588 0.581 14.5s 0.041 0.021 0.029 0.010 0.651 0.677 4.2s
VGGT [25] 0.019 0.008 0.027 0.010 0.611 0.628 76.7s 0.020 0.008 0.027 0.010 0.623 0.641 8.7s
@@
$\pi^{3}$ [29] OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM
StreamVGGT [33] OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM
Fast3R [30] 0.074 0.036 0.024 0.011 0.658 0.682 68.9s 0.061 0.028 0.031 0.013 0.669 0.712 7.4s
CUT3R [26] 0.346 0.243 0.184 0.090 0.579 0.623 18.3s 0.132 0.064 0.056 0.0106 0.669 0.725 5.7s
VGGT [25] 0.029 0.019 0.018 0.009 0.727 0.728 136.1s 0.016 0.010 0.017 0.009 0.735 0.738 13.9s
@@
Comparisons on the ScanNet-50 Dataset. We begin by evaluating FastVGGT on the ScanNet-50 dataset, reporting reconstruction quality using Chamfer Distance (CD). Experiments are conducted with input sequences of 1000, 500, and 100 images, enabling us to assess performance under varying sequence lengths. The results in Table 2 show that although SOTA methods such as $\pi^{3}$ and StreamVGGT achieve strong performance on short sequences, they fail on long sequences due to memory constraints. Methods like Fast3R [30] and CUT3R [26] can process long sequences efficiently, but their reconstruction quality degrades severely. In contrast, FastVGGT delivers substantial acceleration over baseline VGGT across all settings while preserving reconstruction accuracy. Notably, when processing very long sequences ($e.g.$, 1000 images), FastVGGT not only maintains reconstruction fidelity but also significantly mitigates error accumulation, demonstrating robustness and scalability for large-scale 3D reconstruction.
@@ 4.3 Camera Pose Estimation
We begin by evaluating FastVGGT on the ScanNet-50 dataset, reporting reconstruction quality with Chamfer Distance (CD). Experiments are conducted with input sequences of 1000, 500, and 100 images, allowing us to assess performance across different sequence lengths. As summarized in Table 2, state-of-the-art methods such as $\pi^{3}$ and StreamVGGT achieve strong results on short sequences but fail on long ones due to memory limitations. Methods like Fast3R [30] and CUT3R [26] can handle longer sequences efficiently, yet their reconstruction quality drops significantly. In contrast, FastVGGT achieves substantial acceleration over baseline VGGT across all settings while preserving reconstruction accuracy. Notably, on very long sequences ($e.g.$, 1000 images), FastVGGT not only maintains reconstruction fidelity but also mitigates error accumulation, demonstrating robustness and scalability for large-scale 3D reconstruction.
We further evaluate FastVGGT on the 7 Scenes and NRGBD datasets following the CUT3R protocol. Performance is measured by accuracy (Acc), completeness (Comp), and normal consistency (NC) using long-sequence inputs, where keyframes are sampled every 3 or 10 frames. Results in Table 3 and Table 4 show that FastVGGT sustains the robust performance observed on ScanNet-50 and achieves competitive or superior results across all metrics. These experiments confirm that FastVGGT generalizes well beyond ScanNet-50, providing consistent improvements in reconstruction quality and efficiency across diverse scenes, thereby underscoring its practicality for real-world 3D vision applications.

### v2

@@ 1 Introduction
Inferring the 3D geometric structure of a scene from visual inputs, is critical for enabling machines to understand and interact with the physical world. Recent advances in deep learning have catalyzed a paradigm shift in 3D geometric estimation [wang2024dust3r, zhang2025flare, qu2023sg, wang2025vggt, qu2025deocc], enabling a move from iterative, optimization-based pipelines to end-to-end neural networks that directly infer geometry from raw visual inputs. This transformation is exemplified by large-scale architectures like DUSt3R [wang2024dust3r] and its follow-ups [yang2025fast3r, tang2025mv, leroy2024grounding], which showcase a remarkable capacity to reason about complex geometric relationships across image pairs.
Building upon this line of research, VGGT [wang2025vggt] marks a significant advance. Its transformer-based, feed-forward architecture directly regresses key 3D attributes, including camera parameters, depth maps, and point tracks, to achieve highly stable and accurate reconstructions. While this establishes VGGT as a state-of-the-art framework for 3D scene understanding, its scalability is impeded by two critical bottlenecks. First, the model’s reliance on dense global token interactions across views or frames results in prohibitive computational costs. Although used techniques like Flash-Attention mitigate the memory complexity from $O(n^{2})$ to $O(nd)$, the underlying time complexity remains quadratic at $O(n^{2}d)$. Second, the global attention mechanism, essential for capturing inter-frame relations, is susceptible to error accumulation. As the token space expands with each new frame, minor inaccuracies are amplified, leading to significant prediction drift. Collectively, these limitations restrict VGGT’s applicability in large-scale scenarios and motivate the development of more efficient and scalable architectures.
@@
Motivated by the observation of attention redundancy, we adapt the training-free technique of token merging [bolya2022token, bolya2023token, tran2024accelerating, lee2024video] to enhance VGGT’s inference efficiency. Token merging consolidates redundant representations by partitioning tokens into source (src) and destination (dst) sets and merging each src token into its most similar dst counterpart. While effective in 2D vision tasks, its extension to structures designed for 3D geometry understanding remains underexplored. Unlike 2D settings that process single images, VGGT relies on cross-image correspondences, making direct application of token merging highly challenging. To address this, we introduce FastVGGT, a novel, training-free framework that strategically applies token merging to mitigate the Global Attention bottleneck. Our approach begins by preserving the foundational coordinate system. Specifically, tokens from the initial frame, which serves as the global reference for the entire scene, are designated as high-priority destination (dst) tokens and are exempt from being merged to ensure reconstruction stability. Furthermore, to maintain global consistency and preserve fine-grained details, we identify and retain the most salient tokens across all frames, allowing them to bypass the merging process entirely and participate directly in the attention computation. Finally, drawing inspiration from ToMeSD [bolya2023token], we implement region-based random sampling within each subsequent frame. This ensures a spatially balanced selection of src and dst tokens, preventing critical information loss in localized regions during consolidation.
@@ 2.1 Feed-Forward 3D Reconstruction
Building on the foundations of traditional 3D reconstruction [mur2015orb, schonberger2016structure, wang2025look, shen2025evolving, mur2017orb], recent end-to-end learning-based methods leverage neural networks to encode scene priors, substantially enhancing robustness and cross-dataset generalization. Early progress was marked by DUSt3R [wang2024dust3r], which directly regresses view-consistent 3D point maps from only two RGB images without requiring camera calibration. Its successor, MASt3R [leroy2024grounding], introduces confidence-weighted losses to approximate metric scale, further improving reconstruction quality. The current state-of-the-art, VGGT [wang2025vggt] scales this philosophy to a 1.2B-parameter transformer that jointly predicts camera intrinsics, extrinsics, dense depth, point maps, and 2D tracks. However, as input sequences grow longer, the global attention mechanism must capture inter-frame relations within an expanding token space. This not only increases computational overhead but also amplifies noise propagation, making long-sequence predictions more prone to drift. VGGT-Long [deng2025vggt] addresses the drifting issue by aligning sub-maps to suppress error accumulation, but at the cost of significantly reduced inference speed, undermining the efficiency of feed-forward 3D reconstruction. To overcome these challenges, we propose FastVGGT, which accelerates inference and mitigates error accumulation by reducing the number of tokens processed in Global Attention, thereby achieving a balance between efficiency and accuracy.
@@ 2.2 Token Merging
Visual token merging [bolya2022token, renggli2022learning, zeng2022not, haurum2023tokens] was first introduced as a training-free approach to improve the throughput of Vision Transformers (ViTs) [dosovitskiy2020image]. It was later extended to reduce computational cost in tasks such as diffusion, video and language understanding, and video editing [cao2023pumer, choi2024vid, shen2024tempme]. The method partitions tokens into two sets, matches each token with its most similar counterpart, merges them via average pooling, and concatenates the results. Although simple and effective, most applications remain limited to the image domain, while long-form video remains underexplored despite its spatiotemporal tokens being both redundant and interdependent. In parallel, other approaches to token reduction have emerged. TokenLearner [ryoo2021tokenlearner] adaptively selects informative tokens with an MLP. PuMer [cao2023pumer] integrates pruning and merging for vision–language models. Pooling-based methods accelerate attention by averaging token embeddings. While sharing the goal of efficient transformers, our work targets the unique demands of feed-forward 3D reconstruction, where tokens must retain spatial precision and temporal coherence. To this end, we propose FastVGGT, which adapts token merging to VGGT, achieving substantial acceleration while maintaining reconstruction fidelity.
@@ 3.1 Visualization Findings
The phenomenon of attention similarity, often referred to as feature degradation, has been consistently reported in the DINO series [oquab2023dinov2, simeoni2025dinov3, caron2021emerging]. In DINO, while the CLS token becomes increasingly discriminative, patch-level features gradually lose local consistency as they converge toward the CLS token, which undermines performance on dense prediction tasks. A similar limitation appears in VGGT: its global self-attention layers aggregate information across all tokens via weighted averaging, and without explicit regularization or task-specific constraints, the representational space is repeatedly compressed. This compression drives tokens to collapse along a dominant direction, eroding their distinctiveness. At a fundamental level, both DINO and VGGT lack explicit mechanisms to preserve token diversity, leading to a progressive loss of local variation.
It is important to note that the consequences of feature degradation differ markedly between DINO and VGGT. In DINO, the network must handle not only image classification but also dense prediction tasks such as segmentation, where local feature integrity is critical. As training progresses, patch tokens collapse toward the CLS token, weakening dense feature discrimination and directly impairing downstream performance. This motivates corrective strategies such as Gram Anchoring [simeoni2025dinov3]. By contrast, VGGT follows a two-stage design: Global Attention and Frame Attention. Global Attention enforces global consistency by capturing holistic spatial–temporal relationships. Here, the apparent degradation—tokens collapsing toward a dominant subspace—can be interpreted not as a flaw but as deliberate distillation of global semantics. Frame Attention then reintroduces local variability, striking a balance between global abstraction and local differentiation. Moreover, the strong feature similarity observed in Global Attention exposes redundancy that can be exploited to reduce computational cost without compromising performance, providing a natural path to address VGGT’s speed bottleneck.
@@
Figure 4: Overview of the proposed token merging strategy. The pipeline begins with tokenization of input frames. To alleviate the Global Attention bottleneck, we design a five-step procedure: (1) fix initial-frame tokens as destination (Dst) tokens, serving as the global reference for spatial consistency, (2) retain the top-k salient tokens to strengthen correspondences, (3) apply region-based random sampling for balanced selection, (4) fuse source (Src) tokens into their nearest Dst tokens during attention, and (5) decouple tokens via unmerging for dense reconstruction. G-Attn and F-Attn denote Global Attention and Frame Attention, respectively.
@@ 3.2 Token Partitioning
Uniform Token Sampling. Finally, recognizing the dense prediction nature of 3D reconstruction, we ensure uniform intra-frame sampling to avoid local over-compression or redundancy. To this end, we assign dst and src tokens within each frame using a region-based random sampling strategy inspired by the success of ToMeSD [bolya2023token] in diffusion models. Concretely, we first partition the input tokens by frame and arrange them into a 2D grid of image patches. Within each grid cell, dst tokens are sampled according to a predefined merging ratio with stride $K$, while the remaining tokens are designated as src tokens. This region-based strategy ensures spatially balanced merging and prevents artifacts such as the disappearance of large areas. Consequently, the merging process is more coherent, and the reconstructed scene better preserves global structural stability.
@@ 3.4 Token Unmerging Procedure
Dense 3D reconstruction requires per-token outputs (e.g., depth predictions). To satisfy this requirement, we adopt an unmerging operation, inspired by ToMeSD [bolya2023token], which restores the original token resolution and maintains full compatibility with the VGGT architecture. Specifically, suppose two tokens $x_{1},x_{2}\in\mathbb{R}^{c}$ are merged into a single representation:
@@ 4.1 Experimental Setup
We evaluate FastVGGT on three benchmark datasets: ScanNet, NRGBD, and 7 Scenes. For ScanNet, which contains 1,500 scenes, we uniformly sample 50 scenes to form a reproducible benchmark, denoted as ScanNet-50. Our experiments focus on two tasks: camera pose estimation and point map reconstruction. Across both tasks, FastVGGT consistently achieves substantial speedups while maintaining accuracy. The overall architecture follows VGGT, comprising L=24 frame and global attention layers, with Flash-Attention2 [dao2023flashattention] integrated to further accelerate inference. All experiments are conducted on a workstation with an NVIDIA A800 GPU (80 GiB VRAM).
@@
$\pi^{3}$ [wang2025pi] OOM OOM OOM OOM OOM OOM OOM OOM
StreamVGGT[zhuo2025streaming] OOM OOM OOM OOM OOM OOM OOM OOM
Fast3R [yang2025fast3r] 0.684 397.8s 0.701 97.3s 0.711 34.9s 0.723 4.8s
CUT3R [wang2025continuous] 0.786 34.8s 0.774 18.8s 0.775 11.1s 0.767 3.6s
VGGT [wang2025vggt] 0.471 724.6s 0.420 177.5s 0.416 131.4s 0.423 9.1s
@@
$\pi^{3}$ [wang2025pi] OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM
StreamVGGT [zhuo2025streaming] OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM
Fast3R [yang2025fast3r] 0.045 0.027 0.047 0.010 0.616 0.627 43.7s 0.040 0.021 0.056 0.013 0.639 0.657 5.5s
CUT3R [wang2025continuous] 0.179 0.121 0.097 0.043 0.588 0.581 14.5s 0.041 0.021 0.029 0.010 0.651 0.677 4.2s
VGGT [wang2025vggt] 0.019 0.008 0.027 0.010 0.611 0.628 76.7s 0.020 0.008 0.027 0.010 0.623 0.641 8.7s
@@
$\pi^{3}$ [wang2025pi] OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM
StreamVGGT [zhuo2025streaming] OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM OOM
Fast3R [yang2025fast3r] 0.074 0.036 0.024 0.011 0.658 0.682 68.9s 0.061 0.028 0.031 0.013 0.669 0.712 7.4s
CUT3R [wang2025continuous] 0.346 0.243 0.184 0.090 0.579 0.623 18.3s 0.132 0.064 0.056 0.0106 0.669 0.725 5.7s
VGGT [wang2025vggt] 0.029 0.019 0.018 0.009 0.727 0.728 136.1s 0.016 0.010 0.017 0.009 0.735 0.738 13.9s
@@
Comparisons on the ScanNet-50 Dataset. We begin by evaluating FastVGGT on the ScanNet-50 dataset, reporting reconstruction quality using Chamfer Distance (CD). Experiments are conducted with input sequences of 1000, 500, and 100 images, enabling us to assess performance under varying sequence lengths. The results in Table 2 show that although SOTA methods such as $\pi^{3}$ and StreamVGGT achieve strong performance on short sequences, they fail on long sequences due to memory constraints. Methods like Fast3R [yang2025fast3r] and CUT3R [wang2025continuous] can process long sequences efficiently, but their reconstruction quality degrades severely. In contrast, FastVGGT delivers substantial acceleration over baseline VGGT across all settings while preserving reconstruction accuracy. Notably, when processing very long sequences ($e.g.$, 1000 images), FastVGGT not only maintains reconstruction fidelity but also significantly mitigates error accumulation, demonstrating robustness and scalability for large-scale 3D reconstruction.
@@ 4.3 Camera Pose Estimation
We evaluate FastVGGT on the ScanNet-50 dataset for camera pose estimation. Evaluation is conducted using four commonly adopted metrics: Absolute Trajectory Error (ATE), reflecting trajectory-level accuracy; Absolute Rotation Error (ARE), reflecting orientation accuracy; Relative Pose Error in rotation (RPE-rot), and Relative Pose Error in translation (RPE-trans), which jointly characterize local frame-to-frame consistency. As illustrated in Table 5, FastVGGT matches the baseline VGGT on shorter sequences (100 and 300 frames), while on longer sequences (500 and 1000 frames) it substantially reduces pose estimation error. Figure 5 further visualizes pose trajectories, demonstrating FastVGGT’s ability to suppress cumulative drift, thereby underscoring its effectiveness and practicality for real-world deployment.